Solve (p + q)(px + qy) = 1, using Charpit’s method.

1
Expert's answer
2022-02-22T16:21:40-0500

Solution:

The given equation is:

"f(x,y,z,p,q)=(p+q)(px +qy)-1 = p^2x+pqy +pqx+q^2y-1 =0"

Thus the Charpit's auxiliary equations are:

"{dp \over p(p+q)}={dq \over q(p+q)}= {dz \over -p(2px+qy+qx)-q(py+px+2qy)}={dx \over {-(2px+qy+qx)}}={dy \over {-(py+px+2qy)}}"

Taking the first two equations we get "p = aq" for some constant "a" . Putting it in original equation we get:

"q^2 = {1 \over (a+1)(ax+y)}"

and

"p^2 ={a^2 \over (a+1)(ax+y)}" .

Thus "p= {a \over \sqrt{(a+1)(ax+y)}}"

"p= {a \over \sqrt{(a+1)(ax+y)}}" and "q= {1 \over \sqrt{(a+1)(ax+y)}}"

Putting this values of "p, q" in "dz = pdx + qdy" we get:

"dz = {d(ax+y)\over \sqrt{(a+1)(ax+y)}}"

Thus integrating we get,

"z(a+1)^{\frac 12}=2(ax+y)^{\frac 12}+b"

where "a, b" are constants.